write an equation of the line containing the given point and perpendicular to the given line.
(6,-3), 4x+9y=5
type your answer in the form y=mx+b and simplify integer a fraction

Answer by nerdybill(7384)

(6,-3), 4x+9y=5
First, determine slope of:
4x+9y=5
9y=-4x+5
y=(-4/9)x + 5/9
slope is -4/9
.
A line perpendicular has a "negative reciprocal":
(-4/9)m = -1
m = 9/4 (slope of our new line)
.
using m=9/4 and (6,-3) plug into "point-slope" form:
y - y1 = m(x - x1)
y - (-3) = (9/4)(x - 6)
y + 3 = (9/4)x - (9/4)6
y + 3 = (9/4)x - (9/2)3
y + 3 = (9/4)x - 27/2
y = (9/4)x - 27/2 - 3
y = (9/4)x - 27/2 - 6/2
y = (9/4)x - 33/2 (this is what they're looking for)